whether each of these functions is a bijection from R to R. f (x) = x 3 -1 


1
Expert's answer
2021-09-06T14:54:39-0400

Consider the function "f:\R\to\R,\ f(x)=x^3-1." Let us show that this function is an injection. Let "f(x)=f(y)." Then "x^3-1=y^3-1," and hence "x^3=y^3." It follows that "x=y," and "f" is an injective function. Further, let us show that this function is a surjection. Since for any "y\in\R" the equation "f(x)=y" is equivalent to "x^3-1=y," and hence has the solution "x=\sqrt[3]{y+1}\in\R," we conclude that "f" is a surjection. Consequently, the function "f" is a bijection.